Let a = 25 degrees and b = 155 degrees. Which is a + b, in radians?

A. 4π/3

B. π/3

C. 0.8π

D. π
1
Expert's answer
2013-05-20T02:59:54-0400
Let's find a+b in degrees at first: a+b=25+155=180 degrees.
As we know, π is equivalent to 180 degrees, therefore a+b equals π radians. Answer d